# On the potential of quantum walks for modeling financial return distributions

### v1

@@ 1 Introduction
An advantage of using quantum-inspired algorithms in an economic setting is the fact that they may offer practical advantages in their algorithmic implementation. Indeed, quantum computers hold the potential of significantly revolutionizing the handling of computationally challenging problems by outperforming classical computers for selected tasks [15, 16]. This makes quantum computing an attractive area to solve computational problems in finance and economics [17]. It is important to note that a quantum circuit for a one-dimensional quantum walk can be implemented efficiently on a quantum computing device [18].
@@ 2.2 Lévy stable and power law distributions
CDFs of the power law type $\propto x^{-\tilde{\alpha}}$ have been suggested as an alternative way to deal with the occurrence of more extreme events, given their fat-tails like some Lévy stable distributions. Extensive financial data sets have been analysed to investigate the CDFs of the returns defined over different time scales $\Delta t$. Gopikrishnan et al. demonstrated by analyzing the S&P 500 Index (USA), the Nikkei Index (Japan), and the Hang Seng Index (Hong Kong) that for $\Delta t\lesssim 4$ days, the CDF of the returns is consistent with asymptotic power law behavior, where the exponent is given by $\tilde{\alpha}\approx 3$ [25], which is beyond reach for the $0<\alpha\leq 2$ of the Lévy stable distributions [29]. This behavior is often referred to as the inverse cubic law of stock price fluctuations [23, 24], which is applicable to returns defined over time periods varying from 15 seconds to a few days. For longer time scales, a slow convergence to Gaussian behavior is observed [14, 41, 42, 43, 44].
@@
(i) Convergence of the higher moments of Lévy stable and power law distributions is not guaranteed. The $m$-th moment of a power law distribution with exponent $k$ exists for ${m<k-1}$, while all higher moments diverge [47]. This is rather impractical upon analyzing real-life financial data.
@@ 2.3.1 Formalism
where $\xi$, $\theta$ and $\zeta$ are the Caley-Klein parameters [65]. The only relevant parameters for the dynamics of a quantum walker initialized at position $j=0$, are $\eta\equiv\xi-\zeta\in[0,2\pi[$ (proof in A) and $\theta\in[0,\pi[$, since multiplying the $U_{\xi,\theta,\zeta}$ with a global phase does not impact the final result [68]. In many practical applications, one sets $\xi=\zeta=0$, leading to the following coin
@@ 2.3.2 Quantum walk model for financial time series
We now introduce the model for the temporal evolution of the price of financial products. The Wiener process in Eq. (1) is replaced by a quantum walk $Q(t)$
@@
An interesting feature of the quantum walk is that the variance of the resulting position probability distribution $P_{j}(n)$ (Eq. (13)) grows quadratically with the number of time steps $n$. This diffusive behavior is often referred to as “ballistic diffusion” [69, 70]. In a classical random walk, the variance grows linearly with the elapsed time. For now, we have added an additional function $f(t)$ in the stochastic term. This function is intended to account for the ballistic diffusion of the quantum walk, which one might seek to compensate for by setting $f(t)\propto t^{-1/2}$, resulting in more “classical” diffusive behavior. Indeed, the width of the quantum walk probability distribution increases linearly with time $t$, in contrast to GBM, where it scales with $t^{1/2}$. The function $f(t)$ can be chosen such that the standard deviation of the stock returns matches a given value of the stock’s volatility $\sigma$. It is observed that the absolute values of the returns have a positive autocorrelation for time scales ranging from a few minutes to several weeks, whereas the actual values of the returns have non-vanishing but rather insignificant autocorrelations [32, 71, 72]. Based on these observations of so-called volatility clustering [22], the value of $\sigma$ is often kept fixed.
@@ 3.1.1 Dependence on the number of time steps
which leads to a symmetric probability distribution. The choice for this specific initial condition and quantum coin will be clarified in Sections 3.1.2 and 3.1.3. The resulting distributions $P_{j}(n)$ are bipolar, characterized by two distinct peaks attributed to interference effects. The peaks move further away from each other as time progresses [73]. The overall bipolar shape of the probability distribution remains roughly consistent with increasing $n$. The increasing width of the distribution with $n$ can be adjusted for by rescaling the positions relative to the maximal position for a given $n$. A methodology based on a representative number of time steps $n$ can thus capture the relevant features of $P_{j}(n)$.
@@ 3.1.2 Dependence on the initial condition
The level of detail in the distinctive peaks in the probability distribution arising from interference effects in the quantum walk can be mitigated through histogram aggregation, as shown in Fig. 1(c). After this smoothening procedure, one obtains distributions that more closely match those from real-life financial data.
@@ 3.1.3 Dependence on the quantum coin parameters ($\eta=\xi-\zeta$, $\theta$)
The Caley-Klein parameters ($\eta=\xi-\zeta$, $\theta$) in Eq. (8) are a selective instrument to set biases towards either increasing or decreasing asset prices. Fig. 1(d) shows the probability distribution $P_{j}(n=100)$ of three quantum walks with a $U_{\xi,\theta,\zeta}$ quantum coin and the symmetric initial condition (Eq. (15)). The conclusions drawn from the examination of various initial conditions concerning the biases can be extended to this context, where one can introduce an even larger variety of market tendencies by tuning the parameters $\eta$ and $\theta$. By examining the impact of the quantum walk’s initial condition and the Caley-Klein parameters, we can evaluate its effectiveness for modeling the dynamics of extreme market conditions.
@@
where $\kappa_{i}$ is the $i$-th cumulant of the distribution. In Fig. 2, the skewness of the probability distribution $P_{j}(n=100)$ of the quantum walk with the symmetric initial condition of Eq. (15) and the $U_{\xi,\theta,\zeta}$ quantum coin is investigated as a function of $\eta$ and $\theta$. For this symmetric initial condition, $\eta$ and $\theta$ are confined to the interval $[0,\pi/2]$, since all other parameter values yield distributions equivalent to those obtained within the restricted interval (modulo a reflection about $j=0$ in some situations). For these $(\eta,\theta)$-combinations, the skewness adopts negative values, indicating a left-skewed distribution that leans to the right, as can be seen in Fig. 1(d). This can be interpreted as a bias towards positive increments.
@@ 3.2 Decoherence
The decoherence effects simulate the impacts stemming from unpredictable and disrupting information in a supposedly perfectly informed market, and tend to destroy the coherence in the quantum diffusive process. The methodologies designed for introducing decoherence rely on incorporating some sort of randomness into the system (e.g. randomly breaking links with a given probability, introducing a random phase in the quantum coin), which is reminiscent of the random increments in GBM. The introduction of randomness into the decoherence-inducing algorithms aligns with the principles of the efficient market hypothesis.
In the approaches based on the random disruption of links and the introduction of a random phase, there is a tunable parameter that measures the degree to which the quantum walk has collapsed to a classical diffusive process. In Fig. 6, the distribution exhibits fatter tails than the classical case for $p=0.1$ and $p=0.3$, yielding a higher probability of more extreme events than the classical model. This can be useful to overcome issues encountered in classical models lacking this property (e.g. GBM).
@@ 3.3 Advantages over classical methodologies
GBM: inability to capture large price changes great flexibility in shape of return distribution (bipolar, bias towards increasing or decreasing prices, occurrence of extreme changes, inclusion of decoherence effects)
@@
LS&PL: limited insight into the dynamics of the price evolution, only return distributions quantum walk model provides dynamical time evolution, eventually resulting in return distributions
@@ 4 Conclusions
Classical methodologies have been a powerful tool to represent the temporal evolution of asset prices and to quantify stochastic deviations in financial time series. With the prospect of advanced quantum computation technologies, we have explored the potential of the discrete-time quantum walk for modeling financial time series. The concept of a quantum walk adds the flexibility of superposition and interference, which can be used to model competing driving mechanisms in the underlying dynamics. In this respect, quantum walks have the potential to enrich our modeling capabilities beyond those of classical approaches.
@@
An important line for future research lies in the application of the model to historic financial data, in particular to return distributions. This will further deepen our understanding of the financial interpretation of the parameters governing the quantum walk dynamics. In this respect, it can also be further investigated to which exact economic scenarios the model is best applicable.

### v2

@@ 1 Introduction
An advantage of using quantum-inspired algorithms in an economic setting is the fact that they may offer practical advantages in their algorithmic implementation. Indeed, quantum computers hold the potential of significantly revolutionizing the handling of computationally challenging problems by outperforming classical computers for selected tasks [15, 16]. This makes quantum computing an attractive area to solve computational problems in finance and economics [17, 18, 19]. It is important to note that a quantum circuit for a one-dimensional quantum walk can be implemented efficiently on a quantum computing device [20].
@@
The addition of interference effects and quantum entanglement to model asset price evolution through the concept of quantum walks comes at the expense of the introduction of a number of parameters. Throughout this paper, we pay attention to the connection of those parameters to specific market dynamics and provide a systematic study of the parameter space to show their impact on the resulting distributions. An important aspect is to explore whether the advantages of the quantum walk model in capturing the characteristics of financial return distributions justify the added complexity and suggest promising performance from a theoretical perspective. This paper focuses on the theoretical potential of quantum walks to model the temporal evolution of asset prices. A detailed analysis of the model’s performance against real-world asset price returns will be explored in future work.
@@ 2.2 Lévy stable and power law distributions
CDFs of the power law type $\propto x^{-\tilde{\alpha}}$ have been suggested as an alternative way to deal with the occurrence of more extreme events, given their fat-tails like some Lévy stable distributions. Extensive financial data sets have been analyzed to investigate the CDFs of the returns defined over different time scales $\Delta t$. Gopikrishnan et al. demonstrated by analyzing the S&P 500 Index (USA), the Nikkei Index (Japan), and the Hang Seng Index (Hong Kong) that for $\Delta t\lesssim 4$ days, the CDF of the returns is consistent with asymptotic power law behavior, where the exponent is given by $\tilde{\alpha}\approx 3$ [27], which is beyond reach for the $0<\alpha\leq 2$ of the Lévy stable distributions [31]. This behavior is often referred to as the inverse cubic law of stock price fluctuations [25, 26], which is applicable to returns defined over time periods varying from 15 seconds to a few days. For longer time scales, a slow convergence to Gaussian behavior is observed [14, 43, 44, 45, 46].
@@
(i) Convergence of the higher moments of Lévy stable and power law distributions is not guaranteed. The $m$th moment of a power law distribution with exponent $k$ exists for ${m<k-1}$, while all higher moments diverge [49]. This is rather impractical upon analyzing real-life financial data.
@@ 2.3.1 Formalism
where $\xi$, $\theta$ and $\zeta$ are the Cayley-Klein parameters [67]. The only relevant parameters for the dynamics of a quantum walker initialized at position $j=0$, are $\eta\equiv\xi-\zeta\in[0,2\pi[$ (proof in A) and $\theta\in[0,\pi[$, since multiplying the $U_{\xi,\theta,\zeta}$ with a global phase does not impact the final result [70]. In many practical applications, one sets $\xi=\zeta=0$, leading to the following coin
@@ 2.3.2 Quantum walk model for financial time series
We now introduce the model for the temporal evolution of the price of financial products. The central equation of the quantum walk model for financial time series is derived by replacing the Wiener process $W(t)$ in the stochastic differential equation of Eq. (1) by a quantum walk process $Q(t)$
@@
An interesting feature of the quantum walk is that the variance of the resulting position probability distribution $P_{j}(n)$ (Eq. (13)) grows quadratically with the number of time steps $n$. This diffusive behavior is often referred to as “ballistic diffusion” [71, 72]. In a classical random walk, the variance grows linearly with the elapsed time. For now, we have added an additional function $f(t)$ in the stochastic term to control the diffusion properties of the asset price evolution. This function is intended to account for the ballistic diffusion of the quantum walk, which one might seek to compensate for by setting $f(t)\propto t^{-1/2}$, resulting in more “classical” diffusive behavior. Indeed, the width of the quantum walk probability distribution increases linearly with time $t$, in contrast to GBM, where it scales with $t^{1/2}$. Note that this function does not have a quantum mechanical interpretation.
The function $f(t)$ can be chosen such that the standard deviation of the stock returns matches a given value of the stock’s volatility $\sigma$. It is observed that the absolute values of the returns have a positive autocorrelation for time scales ranging from a few minutes to several weeks, whereas the actual values of the returns have non-vanishing but rather insignificant autocorrelations [34, 73, 74]. Based on these observations of so-called volatility clustering [24], the value of $\sigma$ is often kept fixed.
@@ 3.1.1 Dependence on the number of time steps
which leads to a symmetric probability distribution. The choice for this specific initial condition and quantum coin will be clarified in Sections 3.1.2 and 3.1.3. The resulting distributions $P_{j}(n)$ are bimodal, characterized by two distinct peaks attributed to interference effects. The peaks move further away from each other as time progresses [75]. The overall bimodal shape of the probability distribution remains roughly consistent with increasing $n$. The increasing width of the distribution with $n$ can be adjusted for by rescaling the positions relative to the maximal position for a given $n$. A methodology based on a representative number of time steps $n$ can thus capture the relevant features of $P_{j}(n)$.
The distributions in Fig. 1(a) exhibit a bimodal shape, which is rarely observed in short-term financial return distributions (which tend to be unimodal) but can be useful for modeling investor projections [5, 29, 65]. In Sections 3.1.2 and 3.1.3, however, we show the possibility to compute highly asymmetrical return distributions with a substantially less pronounced bimodal shape. Moreover, we will introduce decoherence effects into the quantum walk algorithm in Section 3.2, which also allows one to generate unimodal distributions.
@@ 3.1.2 Dependence on the initial condition
The level of detail in the distinctive peaks in the probability distribution arising from interference effects in the quantum walk can be mitigated through histogram aggregation, as shown in Fig. 1(c). By applying this smoothening procedure, we have removed this spiky behavior that is not expected to be observed in actual financial data.
@@ 3.1.3 Dependence on the quantum coin parameters ($\eta=\xi-\zeta$, $\theta$)
The Cayley-Klein parameters ($\eta=\xi-\zeta$, $\theta$) in Eq. (8) are a selective instrument to set biases towards either increasing or decreasing asset prices. Fig. 1(d) shows the probability distribution $P_{j}(n=100)$ of three quantum walks with a $U_{\xi,\theta,\zeta}$ quantum coin and the symmetric initial condition (Eq. (15)). The conclusions drawn from the examination of various initial conditions concerning the biases can be extended to this context, where one can introduce an even larger variety of market tendencies by tuning the parameters $\eta$ and $\theta$. By examining the impact of the quantum walk’s initial condition and the Cayley-Klein parameters, we can evaluate its effectiveness for modeling the dynamics of extreme market conditions.
@@
where $\kappa_{i}$ is the $i$th cumulant of the distribution. In Fig. 2, the skewness of the probability distribution $P_{j}(n=100)$ of the quantum walk with the symmetric initial condition of Eq. (15) and the $U_{\xi,\theta,\zeta}$ quantum coin is investigated as a function of $\eta$ and $\theta$. For this symmetric initial condition, $\eta$ and $\theta$ are confined to the interval $[0,\pi/2]$, since all other parameter values yield distributions equivalent to those obtained within the restricted interval (modulo a reflection about $j=0$ in some situations). For these $(\eta,\theta)$-combinations, the skewness adopts negative values, indicating a left-skewed distribution that leans to the right, as can be seen in Fig. 1(d). This can be interpreted as a bias towards positive increments.
@@ 3.2 Decoherence
The decoherence effects simulate the impacts stemming from unpredictable and disrupting information in a supposedly perfectly informed market, and tend to destroy the coherence in the quantum diffusive process. The methodologies designed for introducing decoherence rely on incorporating some sort of randomness into the system (e.g. randomly breaking links with a given probability, introducing a random phase in the quantum coin), which is reminiscent of the random increments in GBM.
In the approaches based on the random disruption of links and the introduction of a random phase, there is a tunable parameter that measures the degree to which the quantum walk has collapsed to a classical diffusive process. In Fig. 6, the distributions for $p=0.1$ and $p=0.3$ are non-Gaussian and exhibit fat tails, yielding a higher probability of more extreme events than the classical model. This can be useful to overcome issues encountered in classical models lacking this property (e.g. GBM).
@@ 3.3 Advantages over classical methodologies
GBM: inability to capture large price changes great flexibility in shape of return distribution (bimodal, bias towards increasing or decreasing prices, occurrence of extreme changes, inclusion of decoherence effects)
@@
LS&PL: restricted to morphological fit to return distributions quantum walk model produces return distributions through its time evolution
@@ 4 Conclusions
Classical methodologies have been a powerful tool to represent the temporal evolution of asset prices and to quantify stochastic deviations in financial time series. With the prospect of advanced quantum computation technologies, we have explored the potential of the discrete-time quantum walk for modeling financial time series. The concept of a quantum walk adds the flexibility of superposition and interference, which can be used to model competing driving mechanisms in the underlying dynamics. While the quantum walk model introduces additional parameters, this added complexity is balanced by its ability to model certain financial return characteristics that fall beyond the scope of classical models. In this respect, quantum walks have the potential to enrich our modeling capabilities beyond those of classical approaches.
@@
An important line for future research lies in the application of the model to historic financial data, in particular to return distributions. This will further deepen our understanding of the financial interpretation of the parameters governing the quantum walk dynamics. In this respect, it can also be further investigated to which exact economic scenarios the model is best applicable. Ongoing research includes an in-depth analysis of historical financial data to refine the model’s application and interpretive parameters.
@@
Acknowledgment
This project was conducted with the support of the Special Research Fund of Ghent University (projects BOF/DOC/2023/103 and BOF/BAF/4y/24/1/018).
Data availability
No data was used for the research described in the article.